OpenStudy (anonymous):

Find a polynomial of degree n that has the given zeros. (There are many correct answers.) Zeros x = (–6, –4) Degree n = 2

OpenStudy (rulnick):

(x+6)(x+4) = 0, so x^2+10x+24.

OpenStudy (anonymous):

how do you know where to plug these in? is there a formula?

OpenStudy (rulnick):

If there's a zero at z then x-z = 0.

OpenStudy (rulnick):

x-(-6) = x+6 x-(-4) = x+4